# Assessing treatment efficacy for interval-censored endpoints using multistate semi-Markov models fit to multiple data streams

### v1

@@ 1 Introduction
The primary objective of late-phase vaccine and therapeutic trials is to determine whether an intervention reduces the risk of objective clinical endpoints, symptomatic disease, or death (McLeod and others, 2019). Secondary analyses are often conducted to investigate how the intervention alters the dynamics of clinical progression or to assess the treatment effect on mechanistic endpoints (Food and Drug Administration, 2017). The REGEN-2069 trial found that prophylactic treatment with REGEN-COV, a combination of the monoclonal antibodies (mAb) casirivimab and imdevimab, reduced the risk of symptomatic SARS-CoV-2 infection among individuals at high risk of exposure (O’Brien and others, 2021). Our goals are to assess the PE of REGEN-COV against infection, including asymptomatic infection, to characterize its effect on seroconversion, and estimate the duration of detectable viral shedding.
@@
Multistate models characterize the time-evolution of a biological process through discrete biological states and are natural tools for data assimilation since the model states may be informed by multiple data sources. Observations of continuous observed processes are amenable to analysis with standard parametric and non-parametric methods (Andersen and Keiding, 2002; Cook and Lawless, 2018). Multistate likelihoods for intermittently observed processes involve computationally intractable transition probabilities, except in the case of Markov processes, where transition probabilities can be obtained by solving the Kolmogorov forward equations via a matrix exponential (Kalbfleisch and Lawless, 1985) or by numerical integration (Titman, 2011). A semi-Markov assumption allows transition intensities to depend on the duration of state occupancy and other aspects of the history up to the current state. Misspecification of a model’s operational time scale, e.g., by assuming that a model is time-inhomogeneous Markov rather than semi-Markov, can bias estimates of key functionals, such as sojourn time distributions (Cheung and others, 2022). Unfortunately, transition probabilities can no longer be easily computed for semi-Markov models. One strategy is to approximate a semi-Markov process with a latent continuous-time Markov chain (CTMC) with Coxian phase-type structure that induces occupancy-time dependent sojourns (Titman and Sharples, 2010; Lange and others, 2015). Non-identifiability issues may arise as number of latent states and model parameters increase (Titman and Sharples, 2010). Sampling-based “data augmentation” approaches numerically integrate over the missing paths (Aralis and Brookmeyer, 2019; Barone and Tancredi, 2022), but become computationally expensive and fragile for complex models and when the fraction of missing information is large (Papaspiliopoulos and others, 2007).
We develop a computationally efficient Monte Carlo expectation maximization (MCEM) framework for fitting multistate semi-Markov models when the data are coarsened or subject to measurement error. Our framework is general, rather than model-specific, and explicitly allows for incomplete observation of the state at assessment times and for data to be accrued under continuous and intermittent observation schemes. We incorporate spline-based transition intensities for semi-parametric inference. Our methods allow for a novel and detailed study of the effect of REGN-COV on endpoints related to asymptomatic infection, leading to a more complete understanding of its mechanism and clinical impact.
@@ 2.1 Motivating example: REGEN-COV mAb for COVID-19 prophylaxis
REGEN-2069 was a randomized placebo-controlled trial to evaluate the REGEN-COV mAb as prophylaxis against symptomatic SARS-CoV-2 infection in unvaccinated persons at high risk of infection from a confirmed case within their household (O’Brien and others, 2021). Participants, aged 12 and older with no prior SARS-CoV-2 infection by serology and PCR, were enrolled within 96 hours of a household member’s diagnosis of SARS-CoV-2 infection. Participants were followed continuously for COVID-19 symptoms, which was confirmed by PCR, and were assessed weekly throughout the 28 day EAP for ongoing infection. Serological testing to identify seroconversion following infection was conducted at the end of the EAP.
@@
The data accrue at discrete times and are a coarsened reflection of a participant’s natural history of infection and subsequent clinical and immune response. Table 1(a) illustrates how PCR assessments, symptom history, and serology data combine to enrich the observed history for three hypothetical participants. Participant 001 was PCR+ on day 7, and hence infected in the first week, but did not manifest symptoms or seroconvert. Participant 002 was symptomatically infected. Combining their PCR and symptom history narrows their infection window to day 7, when they were PCR--, to day 12, when their symptoms manifested. Their symptom onset is encoded as occurring in the interval, $(12-\epsilon,12]$. Participant 003 was confirmed asymptomatically infected by final serology. We do not know whether they were infection naïve at their interim assessments since individuals with short shedding durations might escape detection by weekly PCR.
@@
Table 1(b) summarizes the assumptions under which we formulate a model for infection, symptoms, and immunological response. Evidence of SARS-CoV-2 infection may be captured by positive PCR, manifestation of symptoms, or positive serology. Participants who never manifest symptoms or test positive are assumed to be infection naïve at the end of the EAP, i.e., unaffected implies never infected. In this spirit, we treat infection as the occurrence of a participant being measurably affected, as evidenced by viral shedding detectable by nasopharyngeal RT-qPCR. Thus, in our model, infection is the transition from from naïve to PCR+. The other modeling assumptions reflect our understanding of the temporal precedence of viral shedding, symptoms, and seroconversion. For instance, SARS-CoV-2 infected individuals shed detectable virus prior to manifesting symptoms (Puhach and others, 2023). These assumptions inform the formulation of the model states and allowable transitions depicted in Figure 1(c), and identify the model state, or set of possible model states, at each observation time, as shown in the last two columns of Table 1(a).
Observed data Model state
@@
002 7 12-$\epsilon$ + + $\cdot$ Infected in (7, 12-$\epsilon$] 1 2
002 12-$\epsilon$ 12 + + $\cdot$ Symptom onset on D12 2 4
002 12 14 + + $\cdot$ Still PCR+ 4 4
002 14 21 + $\cdot$ PCR+ $\rightarrow$ PCR-- 4 5
@@ 1. Participants who never become symptomatic,
3. PCR-- precludes symptom onset.
4. PCR-- participants cannot again be PCR+.
@@ 2.2 Data for continuously and intermittently observed multistate processes
We proceed to establish notation for representing data from a latent multistate process under a mixture of intermittent and continuous observation schemes. Figures 2(a) and 2(b) provide a visual reference using an illness-death process where disease recurrence is incompletely observed.
@@
(a) $\{Z(\tau),\tau\geq 0\}$, is characterized by states and sojourn times, $(J,W)$. The number of transitions by time $\tau$ is $n(\tau)=\max\{n\geq 0:T_{n}\leq\tau\}$. $T_{n(\tau)}$ is the time at which $Z$ entered $J(\tau)=J_{n(\tau)}$. The time since entry to $J_{3}$ at time $\tau^{\prime}$ is $\tau_{3}=\tau^{\prime}-T_{3}$.
@@
Figure 2: Notation for a latent multistate process (2(a)) and data (2(b)). The true latent path is $Z_{i}=\left\{(H,0),\ (I,W_{1}),\ (H,W_{2}),\ (D,W_{3})\right\}$. We observe $\mathbf{Y}_{i}=\left[\{J(t_{0})=H\},\{J(\tau)=H,\ \tau\in(t_{0},T_{1}]\},\{J(\tau)=I,\ \tau\in(T_{1},t1])\},\ \{J(t_{2})\in(H,I)\},\ \{J(t_{3})=D\}\right]$.
@@ 2.4 Maximum likelihood via Monte Carlo expectation-maximization
[…]
@@ 2.4.1 Data-conditioned Markov proposals for semi-Markov sample paths
The likelihood for a Markov model can be maximized using standard optimization routines to yield a maximum likelihood estimate of $\widetilde{\boldsymbol{\theta}}$, under which we propose sample paths.
@@ 3.1 Analysis of intercurrent events in an illness-death model
We simulated clinical histories for 250 participants in 1,000 hypothetical studies using an illness-death model with Weibull transition intensities with shape parameter greater than 1, so each intensity was 0 at the time of state entry and increased over time. Clinical data was accrued at baseline, which was the time of entry into the healthy state, and at random times spaced roughly every 1 or 3 months over 1 year of follow-up. Confidence intervals for all quantities of interest were computed by summarizing sample paths simulated from the asymptotic distribution of the MLEs. Additional details of the simulation setup and models fit are given in Appendix B.
@@ 3.2 Estimating protective efficacy and natural history in simulated trials
The analysis of the REGEN-2069 trial data in Section 4 aims to characterize the effect of mAb on infection and seroconversion, and to quantify the duration of PCR positivity. Inference is based on the model, depicted in Figure 1(c), which has a naïve uninfected state, and four post-infection states characterized by symptom history and whether SARS-CoV-2 infection is, or has ceased to be, detectable by PCR. A two-stage procedure for calculating the probability of seroconversion given infection is detailed in Section 4.2.1. The data consist of symptom onset times, seroconversion at four weeks, and PCR measurements roughly every week throughout the efficacy assessment period (EAP), with some variation in the visit times to emulate a realistic trial. Trials were simulated from a more complicated semi-Markov model with nine states – a naïve uninfected state, and eight post-infection states characterized by symptom history, PCR positivity, and seroconversion (Table S8). The transitions from naïve to PCR+ without symptoms and from PCR+ without symptoms to diseased were assigned Weibull transition intensities, while the remainder of the transitions had exponential intensities. We set the simulation parameters so that subjects receiving mAb have lower rates of infection and seroconversion along with shorter durations of PCR positivity. The parameterization of the simulation model and ground truths for key quantities of interest are summarized in Tables S9 and S10.
@@ 4.3 Results
Table 4: Model comparison results. The degree and interior knots of the B-Spline intensities are $\kappa$ and $\xi$, respectively. The quantity $l(\theta)-l_{M}(\theta)$ indicates the absolute gain in log-likelihood over the Markov model with exponential transition intensities (larger is better). $\Delta$AIC is computed as $-2(l(\theta)-l_{M}(\theta))+2p$ (smaller is better), where $p$ is the number of model parameters. The log-likelihood of the Markov model is obtained numerically, while those of the semi-Markov model are estimated via Monte Carlo (Appendix A.6).
Figures 3(b) and 3(b) display the pointwise estimates and 95% confidence bands for cumulative incidence of symptomatic infection and all-comer infection, including asymptomatic infection, obtained using the best fit model. The observed incidence of symptomatic infections, which is continuously observed, is fully captured by the confidence bands, an indication that the semi-Markov model is a good fit. We observe a large increase in incidence of infections on day 7 (Figure 3(b)) — these are participants who were infected early, so are pre-/asymptomatic, and who have not yet been identified as infected. Thus the model smoothly estimates a sharp increase in infections shortly after randomziation and avoids the artifactual bump at day 7 that is an artifact of weekly PCR assessment. The discrepancy between the observed and estimated incidence is greatest in the first week, when the majority of infections occurred — 64.5% (95% CI: 57.2%, 72.0%) of the total incidence in the placebo arm and 63.3% (95% CI: 55.6%, 71.3%) in the mAb arm. Furthermore the model has higher cumulative incidence at day 28, which reflects infections of short duration that are missed by intermittent sampling.
@@
Infected participants on the mAb arm were 31.9% (95% CI: 22.3%, 44.6%) as likely to seroconvert following infection compared with participants on the placebo arm (Figure 3(c)). The rate of seroconversion by day 28 among asymptomatically infected placebo participants was lower than among symptomatically infected placebo participants, 64.3% 95% CI: 52.2%, 77.7%) versus 77.5% (95% CI: 65.5%, 88.3%), but significantly higher than the rate of seroconversion among asymptomatic mAb participants, 24.6% (95% CI: 12.2%, 40.6%). This dynamic lends support to the hypothesis that high viral loads increase both the severity of disease and the intensity of the ensuing immune response leading to expression of anti-nucleocapsid antibodies.

### v2

@@ 1 Introduction
The primary objective of late-phase vaccine and therapeutic trials is to determine if an intervention reduces the risk of objective clinical endpoints, symptomatic disease, or death (McLeod and others, 2019). Secondary analyses examine how the intervention alters the dynamics of clinical progression and assess the treatment effect on mechanistic endpoints (Food and Drug Administration, 2017). The REGEN-2069 trial found that prophylactic treatment with REGEN-COV, a combination of the monoclonal antibodies (mAb) casirivimab and imdevimab, reduced the risk of symptomatic SARS-CoV-2 infection among individuals at high risk of exposure (O’Brien and others, 2021). Our goals are to assess the PE of REGEN-COV against infection, including asymptomatic infection, to characterize its effect on seroconversion, and estimate the duration of detectable viral shedding.
@@
Multistate models characterize the time-evolution of a biological process through discrete biological states and are natural tools for data assimilation since the model states may be informed by multiple data sources. Observations of continuous observed processes are amenable to analysis with standard parametric and non-parametric methods (Andersen and Keiding, 2002; Cook and Lawless, 2018). Multistate likelihoods for intermittently observed processes involve computationally intractable transition probabilities except in the case of Markov processes where transition probabilities can be obtained by solving the Kolmogorov forward equations via a matrix exponential (Kalbfleisch and Lawless, 1985) or by numerical integration (Titman, 2011). A semi-Markov assumption allows transition intensities to depend on the duration of state occupancy and other aspects of the history up to the current state. Misspecification of a model’s operational time scale, e.g., by assuming that a model is time-inhomogeneous Markov rather than semi-Markov, can bias estimates of key functionals, such as sojourn time distributions (Cheung and others, 2022). Unfortunately, transition probabilities can no longer be easily computed for semi-Markov models. One strategy is to approximate a semi-Markov process with a latent continuous-time Markov chain (CTMC) with Coxian phase-type structure that induces occupancy-time dependent sojourns (Titman and Sharples, 2010; Lange and others, 2015). Phase-type latent CTMCs may not be identifiabile when the true process is nearly Markovian or the data cannot resolve more complicated temporal dynamics (Titman and Sharples, 2010). Existing sampling-based “data augmentation” approaches numerically integrate over the missing paths (Aralis and Brookmeyer, 2019; Barone and Tancredi, 2022), but become computationally expensive and fragile for complex models and when the fraction of missing information is large (Papaspiliopoulos and others, 2007).
We develop a computationally efficient Monte Carlo expectation maximization (MCEM) framework for fitting multistate semi-Markov models when the data are coarsened or subject to measurement error. Our framework is general, rather than model-specific, and explicitly allows for incomplete observation of the state at assessment times and for data to be accrued under continuous and intermittent observation schemes. We incorporate covariate effects and spline-based transition intensities for semi-parametric inference. Our methods allow for a novel and detailed study of the effect of REGEN-COV on endpoints related to asymptomatic infection, leading to a more complete understanding of its mechanism and clinical impact.
@@ 2.1 Motivating example: REGEN-COV mAb for COVID-19 prophylaxis
REGEN-2069 was a randomized placebo-controlled trial to evaluate the REGEN-COV mAb as prophylaxis against symptomatic SARS-CoV-2 infection in unvaccinated persons at high risk of infection from a confirmed case within their household (O’Brien and others, 2021). Participants, aged 12 and older with no prior SARS-CoV-2 infection by serology and PCR, were enrolled within 96 hours of a household member’s diagnosis of SARS-CoV-2 infection. Participants were monitored for symptoms and SARS-CoV-2 infection was confirmed by PCR at presentation. Participants’ infection status was also assessed weekly by PCR for viral shedding, and serological testing conducted at the end of the EAP to identify seroconversion following infection.
@@
The data accrue at discrete times and are a coarsened reflection of the natural history of infection and subsequent clinical and immune response. Table 1(a) illustrates how PCR assessments, symptom history, and serology data combine to enrich the data for three hypothetical participants. Participant 001 was PCR+ on day 7, and hence infected in the first week, but did not manifest symptoms or seroconvert. Participant 002 was symptomatic and had infection confirmed by PCR at symptom onset, though their infection would not have been detected on the basis of regularly scheduled PCR assessments and final serology. Combining their PCR and symptom history narrows their infection window to day 7, when they were PCR--, to day 9, when their symptoms manifested. Their symptom onset is encoded as occurring in the interval, (9-$\epsilon$, 9] where $\epsilon$ is an arbitrary duration that is short relative to the time-resolution of the data, e.g., 1 hour. Participant 002 stopped shedding detectable virus by day 14. Participant 003 was confirmed asymptomatically infected by final serology. We do not know whether they were infection naïve at their interim assessments since individuals with short shedding durations might escape detection by weekly PCR.
@@
Table 1(b) summarizes the assumptions under which we formulate a biological model for infection, symptoms, and immunological response. Evidence of SARS-CoV-2 infection may be captured by positive PCR, manifestation of symptoms, or positive serology. Participants who never manifest symptoms or test positive are assumed to be infection naïve at the end of the EAP, i.e., unaffected implies never infected. In this spirit, we treat infection as the occurrence of a participant being measurably affected, as evidenced by viral shedding detectable by nasopharyngeal RT-qPCR. Thus, in our model, infection is the transition from from naïve to PCR+. The other modeling assumptions reflect our understanding of the temporal precedence of viral shedding, symptoms, and seroconversion. For instance, SARS-CoV-2 infected individuals shed detectable virus prior to manifesting symptoms (Puhach and others, 2023). These assumptions inform the formulation of the model states and allowable transitions depicted in Figure 1(c), and identify the model state, or set of possible model states, at each observation time, as shown in the last two columns of Table 1(a).
Before proceeding, we highlight several aspects of the data that complicate statistical inference. First, the data intermingle continuous observations of one aspect of the process, symptom onset, with intermittent observations of other aspects of progression, namely, infection and viral shedding. Second, the true state for participants whose infection is confirmed by serology alone is unknown at interim assessment times. Finally, the data do not preclude the occurrence of multiple transitions within an inter-observation interval and there is some variability in assessment times around scheduled assessments, e.g., a PCR assessment scheduled for study day 7 could have been conducted at day 9. In consideration of these complications, the next sections will develop a framework using continuous-time models for the latent biological process, though we note a similar development is possible for multistate processes that evolve discretely in time.
Observed data at end time Model state
@@
002 7 9-$\epsilon$ + + $\cdot$ Infected in (7, 9-$\epsilon$] 1 2
002 9-$\epsilon$ 9 + + $\cdot$ Symptom onset on D9 2 4
002 9 14 + $\cdot$ PCR+ $\rightarrow$ PCR-- in (9, 14] 4 4
002 14 21 + $\cdot$ No change in PCR– 4 5
@@ 1. Participants who never become symptomatic,
3. Symptom onset precedes PCR+ $\longrightarrow$ PCR--.
4. No transitions from PCR-- $\longrightarrow$ PCR+.
@@ 2.2 Data for continuously and intermittently observed multistate processes
We proceed to establish notation for representing data from a latent multistate process under a mixture of intermittent and continuous observation schemes while allowing for incomplete observation of the true state. Figures 2(a) and 2(b) provide a visual reference using a progressive illness-death process where disease recurrence is incompletely observed.
@@
(a) $\{Z(\tau),\tau\geq 0\}$, is characterized by states and sojourn times, $(J,W)$. The number of transitions by time $\tau$ is $n(\tau)=\max\{n\geq 0:T_{n}\leq\tau\}$. $T_{n(\tau)}$ is the time at which $Z$ entered $J(\tau)=J_{n(\tau)}$. The time since entry to $J_{2}$ at time $\tau^{\prime}$ is $\tau_{2}=\tau^{\prime}-T_{2}$.
@@
Figure 2: Notation for a latent multistate process (2(a)) and data (2(b)). The true path, $\left\{(H,0),\ (I,W_{1}),\ (D,W_{2})\right\}$, is continuously observed from $t_{0}$ until $t_{1}$. The participant is observed to be alive at $t_{2}$ and the exact time of death is recorded at $t_{3}$. Hence, we observe [$\left[\{J(t_{0})=H\},\ \{J(\tau)=H,\ \tau\in(t_{0},t_{1}]\},\{J(t_{2})\in(H,I)\},\ \{J(t_{3})=D\}\right]$.
@@ 2.4 Maximum likelihood via Monte Carlo expectation-maximization
where $\bar{\nu}_{i,m}=\nu_{i,m}/\sum_{j=1}^{M_{i}}\nu_{i,j}$ are the self-normalized importance weights (Liu and Liu, 2001). Early iterations of the MCEM algorithm will typically yield large increases in the Q-function. Hence, setting the initial ESS at some modest target, e.g., $M_{i}^{\star}=10$ to $25$, avoids wasteful computation when the algorithm is far from the MLE. We suggest a larger starting ESS, e.g., $M_{i}^{\star}=50$ to $100$, for complicated models to help the algorithm avoid local modes. The ascent MCEM algorithm halts iteration of the E- and M-steps when the upper limit of a one-sided $1-\gamma$ confidence interval (CI) for the change in the Q-function rules out a preset tolerance.
@@ 2.4.1 Data-conditioned Markov proposals for semi-Markov sample paths
The likelihood for a Markov model can be maximized using standard optimization routines to yield a maximum likelihood estimate (MLE) of $\widetilde{\boldsymbol{\theta}}$, under which we propose sample paths.
@@ 3.1 Analysis of intercurrent events in an illness-death model
We simulated clinical histories for 250 participants in 1,000 hypothetical studies using an illness-death model with Weibull transition intensities with shape parameter greater than 1, so each intensity was 0 at the time of state entry and increased over time (Table S5). Clinical data was accrued at baseline, which was the time of entry into the healthy state, and at random times spaced roughly every 1 or 3 months over 1 year of follow-up. Confidence intervals for all quantities of interest were computed by summarizing sample paths simulated from the asymptotic distribution of the MLEs. Additional details of the simulation setup and models fit are given in Appendix B.
@@ 3.2 Estimating protective efficacy and natural history in simulated trials
The analysis of the REGEN-2069 trial data in Section 4 aims to characterize the effect of mAb on infection and seroconversion, and to quantify the duration of PCR positivity. Inference is based on the model, depicted in Figure 1(c), which has a naïve uninfected state, and four post-infection states characterized by symptom history and whether SARS-CoV-2 infection is, or has ceased to be, detectable by PCR. A two-stage procedure for calculating the probability of seroconversion given infection is detailed in Section 4.2.1. The data consist of symptom onset times, seroconversion at four weeks, and PCR measurements roughly every week throughout the efficacy assessment period (EAP), with some variation in the visit times to emulate a realistic trial. Trials were simulated from a more complicated semi-Markov model with nine states – a naïve uninfected state, and eight post-infection states characterized by symptom history, PCR positivity, and seroconversion (Table S8). The transitions from naïve to PCR+ without symptoms and from PCR+ without symptoms to diseased were assigned Weibull transition intensities, while the remainder of the transitions had exponential intensities. The shape parameter for the infection transition, $1\rightarrow 2$, was less than 1 to reflect a decreasing force of infection, and the shape parameter for symptom onset was greater than 1 to induce a delay in symptom onset. We set the simulation parameters so that subjects receiving mAb have lower rates of infection and seroconversion along with shorter durations of PCR positivity. The parameterization of the simulation model and ground truths for key quantities of interest are summarized in Tables S9 and S10.
@@ 4.3 Results
Table 4: Model comparison results. The degree and interior knots of the B-Spline intensities are $\kappa$ and $\xi$, respectively. The quantity $\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y})$ indicates the absolute gain in log-likelihood over the Markov model with exponential transition intensities (larger is better). $\Delta$AIC is computed as $-2(\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y}))+2p$ (smaller is better), where $p$ is the number of model parameters. The log-likelihood of the Markov model is obtained numerically, while those of the semi-Markov model are estimated via Monte Carlo (Appendix A.6).
Figures 3(b) and 3(b) display the pointwise estimates and 95% confidence bands for cumulative incidence of symptomatic infection and all-comer infection, including asymptomatic infection, obtained using the best fit model. The observed incidence of symptomatic infections, which is continuously observed, is fully captured by the confidence bands, an indication that the semi-Markov model is a good fit. We observe a large increase in incidence of infections on day 7 (Figure 3(b)) — these are participants who were infected early, so are pre-/asymptomatic, and who have not yet been identified as infected. Thus, the model smoothly estimates a sharp increase in infections shortly after randomziation and avoids the artifactual bump at day 7 that is an artifact of weekly PCR assessment. The discrepancy between the observed and estimated incidence is greatest in the first week, when the majority of infections occurred — 64.5% (95% CI: 57.2%, 72.0%) of the total incidence in the placebo arm and 63.3% (95% CI: 55.6%, 71.3%) in the mAb arm. Furthermore the model has higher cumulative incidence at day 28, which reflects infections of short duration that are missed by intermittent sampling.
@@
Infected participants on the mAb arm were 31.9% (95% CI: 22.3%, 44.6%) as likely to seroconvert following infection compared with participants on the placebo arm (Figure 3(c)). The rate of seroconversion by day 28 among asymptomatically infected placebo participants was lower than among symptomatically infected placebo participants, 64.3% (95% CI: 52.2%, 77.7%) versus 77.5% (95% CI: 65.5%, 88.3%), but significantly higher than the rate of seroconversion among asymptomatic mAb participants, 24.6% (95% CI: 12.2%, 40.6%). This dynamic lends support to the hypothesis that high viral loads increase both the severity of disease and the intensity of the ensuing immune response leading to expression of anti-nucleocapsid antibodies.